OpenStudy (anonymous):

Which is equal to 5^4 • 5^8? A. 5^12 B. 5^4 C. 5^2 D. 5^−4

Nnesha (nnesha):

when we multiply same bases we should add their exponents \[\huge\rm x^m \times x^n = x^{m+n}\]

OpenStudy (anonymous):

help no straight answer plz

OpenStudy (anonymous):

ok

Nnesha (nnesha):

don't worry :-)

Nnesha (nnesha):

so 5^4 times 5^8 = ???

OpenStudy (anonymous):

244140625

OpenStudy (anonymous):

???

Nnesha (nnesha):

...nope we don't need calculator... just add exponents exponent rule! \[\huge\rm \color{green}{x}^\color{red}{m} \times \color{green}{x}^\color{blue}{n} = \color{green}{x}^{\color{red}{m}+\color{blue}{n}}\]

OpenStudy (anonymous):

oh ok so its a right???

OpenStudy (anikate):

its easier than you think lol

Nnesha (nnesha):

yep that's right \[\huge\rm \color{green}{5}^\color{red}{4} \times \color{green}{5}^\color{blue}{8} = \color{green}{5}^{\color{red}{4}+\color{blue}{8}}\]

OpenStudy (anonymous):

wow thanks you helped me

Nnesha (nnesha):

np :-)

OpenStudy (anonymous):

d:P